eviant:

The median age for a first marriage in the United States for women was 25.9 in 2009 and 26.1 in 2010. Use an exponential model to predict the median age for women in 2019, where x is the number of years since 2009.

eviant:

@Zepdrix

Zepdrix:

So if we let 2009 be our starting year, \(\large\rm f(x)=a\cdot b^x\) Then x=0 since 2009 to start, (0, 25.9) \(\large\rm f(0)=25.9=a\cdot b^0\) So that gives us our a value, a=25.9 \(\large\rm f(x)=25.9\cdot b^x\)

Zepdrix:

In 2010, our x will be 1, (1, 26.1) \(\large\rm f(1)=26.1=25.9\cdot b^1\) So we can find our b value using some division, \(\large\rm b=\dfrac{26.1}{25.9}\approx 1.007722\)

Zepdrix:

\(\large\rm f(x)=25.9(1.007722)^x\)

Zepdrix:

And then in 2019, our x will be 10, \(\large\rm f(10)=25.9(1.007722)^{10}\)

Zepdrix:

Let your calculator do the rest :D